OpenStudy (anonymous):

Look at the following sequence... 5, 10.5, 22.05, 46.305 If it is a geometric sequence choose the common ratio, if not then choose not geometric

OpenStudy (anonymous):

@surjithayer

OpenStudy (campbell_st):

you need to find the common ratio so compare the ratio of terms does \[\frac{a_{2}}{a_{1}} = \frac{a_{3}}{a_{2}} = \frac{a_{4}}{a_{3}}=...\]

OpenStudy (campbell_st):

so if you keep getting the same value, then that is the common ratio...

OpenStudy (anonymous):

Which is which? Which is a^2,

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

so the sequence is \[a_{1} ~1st~ term, ~~a_{2} ~ 2nd~term, ~~a_{3} ~ 3rd ~term...... \]

OpenStudy (campbell_st):

if the ratio of terms is the same, then it is geometric

OpenStudy (anonymous):

I got 2.1. So it's a common ratio? @campbell_st